# WaterSeeker: Pioneering Efficient Detection of Watermarked Segments in Large Documents

### v5

@@ 1 Introduction
As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Megías et al., 2022; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023). LLM watermarking technology has emerged to tackle these issues by embedding specific information (watermarks) during text generation, allowing for accurate detection through specialized algorithms. Current watermark detection methods (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024; Aaronson and Kirchner, 2022; Lu et al., 2024; Lee et al., 2023; Hu et al., 2024; Wu et al., 2023) first calculate watermark scores for individual tokens through single token detection, then compute statistics across the entire document for classification. While these full-text detection methods effectively distinguish between fully watermarked and non-watermarked texts, they fail in real-world scenarios where LLMs generate only brief segments within longer documents, due to dilution effects, as shown in Figure 1. To the best of our knowledge, the WinMax algorithm (Kirchenbauer et al., 2024) is the only work addressing this limitation by examining all possible window sizes and selecting the maximum statistical score across all windows as the final detection result, but suffers from high time complexity.
@@ 2 Related Work
Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023; Pan et al., 2024). The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024; Wu et al., 2023) categorizes vocabulary into green and red lists, biasing towards green tokens during generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For these methods, single token detection determines whether each token belongs to the green list, while full-text detection involves calculating the z-score of green tokens across the entire document; exceeding a threshold indicates watermarking.

### v6

@@ 1 Introduction
As large language models (LLMs) generate high-quality text, they address practical challenges but also raise concerns such as misinformation (Megías et al., 2022; Chen and Shu, 2024) and copyright infringement (Rillig et al., 2023). LLM watermarking technology has emerged to tackle these issues by embedding specific information (watermarks) during text generation, allowing for accurate detection through specialized algorithms. Current watermark detection methods (Kirchenbauer et al., 2023; Zhao et al., 2024; Liu et al., 2024b; Aaronson and Kirchner, 2022; Lu et al., 2024; Lee et al., 2023; Hu et al., 2024; Wu et al., 2023) first calculate watermark scores for individual tokens through single token detection, then compute statistics across the entire document for classification. While these full-text detection methods effectively distinguish between fully watermarked and non-watermarked texts, they fail in real-world scenarios where LLMs generate only brief segments within longer documents, due to dilution effects, as shown in Figure 1. To the best of our knowledge, the WinMax algorithm (Kirchenbauer et al., 2024) is the only work addressing this limitation by examining all possible window sizes and selecting the maximum statistical score across all windows as the final detection result, but suffers from high time complexity.
@@ 2 Related Work
Currently, mainstream LLM watermarking methods involve modifying the inference phase by altering logits or influencing token sampling (Liu et al., 2023; Pan et al., 2024; Liu et al., 2024c). The KGW family (Kirchenbauer et al., 2023; Zhao et al., 2024; Hu et al., 2024; Liu et al., 2024b; Wu et al., 2023; Liu et al., 2024a; He et al., 2024; Huo et al., 2024) categorizes vocabulary into green and red lists, biasing towards green tokens during generation. The bias value is typically determined by the parameter $\delta$, which reflects the watermark strength. For these methods, single token detection determines whether each token belongs to the green list, while full-text detection involves calculating the z-score of green tokens across the entire document; exceeding a threshold indicates watermarking.
@@ 7 Conclusion
Acknowledgments
This work is primarily supported by the Key Research and Development Program of China (No. 2024YFB3309702). We would like to express our gratitude to the anonymous ARR October reviewers (Reviewer 6E27, 1X1q, B2fg) and Area Chair zwno for their valuable feedback and suggestions that helped improve this paper. Additionally, we extend our sincere thanks to Hanyu Xue for his assistance with mathematical derivations, which provided valuable insights throughout this research.